OpenStudy (anonymous):

Find all solutions in the interval [0, 2π). 4 sin2 x - 4 sin x + 1 = 0

OpenStudy (callisto):

Do you mean\[4 sin^2 x - 4 sin x + 1 = 0\]?

OpenStudy (anonymous):

yes sorry

OpenStudy (callisto):

Nope :) Treat sinx as a, you'll get 4a^2 - 4a +1 =0 Can you solve this equation?

OpenStudy (anonymous):

not really

OpenStudy (callisto):

Can you factorize the left side?

OpenStudy (anonymous):

no

OpenStudy (callisto):

Hint: use \(a^2 -2ab +b^2 = (a-b)^2\)